# Deciphering the Milky Way disc formation time encrypted in the chrono-kinematics of the bar

### v1

@@ Abstract
pubyear: 2024pagerange: Deciphering the Milky Way disc formation time encrypted in the chrono-kinematics of the barA
@@ 1 Introduction
A disc is a typical component of many galaxies in the Universe. Disc formation and evolution in galaxies are studied with high-redshift observations of morphology (e.g. Nelson et al., 2023; Kartaltepe et al., 2023; Robertson et al., 2023) and gas kinematics (e.g. Wisnioski et al., 2019; Übler et al., 2019; de Graaff et al., 2024), as well as with cosmological simulations (e.g. Stern et al., 2021; Hopkins et al., 2023; Semenov et al., 2024; Dillamore et al., 2024). In the Milky Way, because we can resolve individual stars and measure their complete 6D phase space positions, our Galaxy is an excellent laboratory to study disc formation and evolution (e.g. Carollo et al., 2019; Sestito et al., 2020; Di Matteo et al., 2020; Chandra et al., 2023; Zhang et al., 2024b; Nepal et al., 2024; Gallart et al., 2024).
@@
The formation time of the Milky Way disc can be studied both using metallicity as a clock of Galactic evolution (Norris et al., 1985; Sestito et al., 2020; Di Matteo et al., 2020; Belokurov & Kravtsov, 2022; Chandra et al., 2023; Zhang et al., 2024b; Fernández-Alvar et al., 2024; Viswanathan et al., 2024) as well as using stellar ages directly (Miglio et al., 2021; Rix et al., 2022; Xiang & Rix, 2022; Queiroz et al., 2023; Nepal et al., 2024). The metallicity distribution of the Galactic disc has received much attention in the past decades, especially at the metal-poor regime, which is potentially linked to the epochs of its early formation. Observations invoking the existence of a metal-weak thick disc are discussed periodically (e.g. Norris et al., 1985; Morrison et al., 1990; Chiba & Beers, 2000; Carollo et al., 2010, 2019; An & Beers, 2020; Di Matteo et al., 2020).
@@
Recently, very metal-poor (VMP) stars, which are expected to be the oldest stars in the Milky Way, with disc-like orbits discovered by Sestito et al. (2019, 2020) have revived the discussion. Although the disc-like VMP stars could have multiple origins  (see Sestito et al., 2021), the hypothesis that they come from an early proto-disc has attracted the most attention (e.g. Di Matteo et al., 2020; Mardini et al., 2022; Bellazzini et al., 2024). Using APOGEE data, Belokurov & Kravtsov (2022) created a pure sample of in-situ MW stars and revealed that the net azimuthal velocity of the Galaxy increases rapidly across a short metallicity range around $\mathrm{[M/H]}\sim-1.3$. Comparing to a variety of numerical models of galaxy formation, Belokurov & Kravtsov (2022) conclude that the fast observed spin-up of the Milky Way marks the emergence of the stable and dominant Galactic stellar disc. Note that while the metallicity scale of the Galactic spin-up is not very low, the corresponding lookback time is impressively large, corresponding to redshifts above $z>3$. Further support for this hypothesis has been found in independent studies of other simulation suites (e.g. Semenov et al., 2024; Dillamore et al., 2024; Chandra et al., 2023). Indeed, as these studies demostrate, our Galaxy was likely one of the first to form a stable stellar disc amongst objects of similar halo mass. With metallicities from the Gaia BP/RP (XP) spectra (e.g. Andrae et al., 2023), we can access an unprecedentedly large sample of metal-poor stars. Recent analyses of such Gaia XP-based samples bolster the idea that the Milky Way disc formed around $\mathrm{[M/H]}\sim-1.5$ to $-1$ (Chandra et al., 2023; Zhang et al., 2024b). Other origins for the VMP stars on disc-like orbits have also been discussed. These can form in the proto-Galaxy  (Sestito et al., 2021; Horta et al., 2024; Zhang et al., 2024b), or can be trapped by a decelerating bar (Dillamore et al., 2023; Li et al., 2023; Yuan et al., 2023).
@@
Although the stellar metallicity is usually measured more accurate than the age, the metallicity evolution also depends on the star formation history (sometimes leading to non-monotonic age-metallicity relation), which complicates the interpretation of results. Accurate age measurements of disc stars can directly pin down the formation epoch of the Milky Way disc, but this is hard, and very few robust age estimates exist for low-metallicity stars. Miglio et al. (2021) assessed the asteroseismological age of stars with light curves measured by the Kepler satellite and found the mean age of stars in the thick disc after deconvolving the age uncertainty is consistent with $\sim 11$ Gyr. Similarly, Queiroz et al. (2023) used the stellar ages from isochrone fitting and found the same age for the thick disc. Recently, Nepal et al. (2024) found disc populations of stars older than $13$ Gyr using stellar parameters from the Gaia RVS+XP (Guiglion et al., 2024), which hints that the Galactic disc may have formed within $1$ Gyr after the Big Bang. However, reconciling the early disc formation scenario with the cosmological simulations is challenging. In numerical models, the morphology of a galaxy appears to change quickly from spheroidal to disc-like when the galaxy reaches a critical mass threshold (Hopkins et al., 2023; Dillamore et al., 2024). Taking the diversity of the assembly histories into account, it appears unusual for a Milky Way-like galaxy to reach such a mass threshold much earlier than redshift $\sim 3$ (Belokurov & Kravtsov, 2022).
@@
In this work, we provide a new method to measure the disc formation epoch that uses the kinematics of the inner Galaxy, which is independent of previous measurements. The Galactic bar dominates the inner Galaxy and is identified from the stellar density distribution at the centre (Nakada et al., 1991; Stanek et al., 1997; Wegg & Gerhard, 2013). Other than the spatial overdensity caused by the Galactic bar, it also induces distinct kinematic features in the inner Galaxy, which are useful when tracing the bar signatures. Due to the streaming motion of bar-supporting stars and their elongated orbits, the kinematics of the Galactic bar is characterised by the quadrupole radial velocity pattern (Bovy et al., 2019; Queiroz et al., 2021; Hey et al., 2023; Vislosky et al., 2024; Liao et al., 2024; Zhang et al., 2024a) and the bi-symmetric pattern in the fractional radial velocity map (Zhang et al., 2024a).
@@
The structure and formation of the bar has been explored intensively with N-body simulations (e.g. Athanassoula & Misiriotis, 2002; Fragkoudi et al., 2017; Fujii et al., 2018; Ghosh et al., 2024), and recently, with the Auriga cosmological simulation suite (Grand et al., 2017; Fragkoudi et al., 2020; Merrow et al., 2024; Fragkoudi et al., 2024). Several factors can impact the bar formation in a galaxy, such as disc kinematics (e.g. Toomre, 1964; Athanassoula & Sellwood, 1986), disc mass fraction (e.g. Fujii et al., 2018), and gas fraction (e.g. Athanassoula et al., 2013). Many studies used N-body simulations to investigate the response of different stellar populations to the Galactic bar (e.g. classical bulge: Saha et al. 2012, 2016, the stellar halo: Pérez-Villegas et al. 2017, and the Galactic disc: Debattista et al. 2017; Fragkoudi et al. 2017). Debattista et al. (2017) demonstrated that cold populations form a strong, thin, peanut-shaped bar, while hot populations form a weak, thick, boxy-shaped bar. As kinematically colder population respond to the Galactic bar more efficiently, the cold galactic disc provides the stars for the bar formation (Toomre, 1964; Fragkoudi et al., 2017; Debattista et al., 2017; Boin et al., 2024). Moreover, early interactions with large satellites can seed bar formation. For example, Merrow et al. (2024) show that an accretion event similar to the GS/E merger (Belokurov et al., 2018; Helmi et al., 2018) in the MW can be responsible for the bar formation, as long as the GS/E progenitor is not too massive.
@@
Long-period variables (LPV) are bright stars that follow tight period-luminosity relations (Rau et al., 2019; Sanders, 2023) and hence, are often used to trace the structure and kinematics of the inner Galaxy (e.g. Catchpole et al., 2016; Grady et al., 2020; Sanders et al., 2022; Hey et al., 2023; Zhang et al., 2024a). Mira variables, a subcategory of LPVs, are high-amplitude thermally pulsating asymptotic giant branch stars. Their distances can be calibrated via the period-luminosity relation of O-rich Mira variables in the Milky Way (Sanders, 2023) and can have astrometry from Gaia DR3. Mira variables also exhibit a period-age relation: Miras with shorter periods have older ages (Feast & Whitelock, 2000; Grady et al., 2019; Trabucchi & Mowlavi, 2022; Zhang & Sanders, 2023). Zhang & Sanders (2023) re-calibrated the period-age relation of O-rich Mira variables to ensure consistency with the ages of Miras in globular clusters. Mira variables have been used to trace the chronological evolution of the Milky Way structures (e.g. Catchpole et al., 2016; Grady et al., 2020), and recently, Sanders et al. (2024) used the Mira variables in the nuclear stellar disc to constrain the bar formation epoch of the Milky Way.
@@ 2.1 O-rich Mira variable selection
Although Mira variables with a period shorter than $150$ days exist, we exclude LPV candidates with $\mathrm{period}<150$ days to remove potential short-period(SP)-red stars contamination. (e.g., Feast & Whitelock, 2000). Kinematically, the SP-red stars behave similarly to longer-period Mira variables, which have a lower velocity dispersion. This indicates that rather than being older these stars have smaller ages and gave sustained less heating. A related result is found in Zhang & Sanders (2023), in which the kinematics of the Mira sample with a period of 80 to 150 days is kinematically colder than expected. We omit the contaminating stars, many of which are likely SP-red objects from our analysis by neglecting Mira candidates with a period shorter than 150 days, which also removes the oldest Mira candidates ($\mathrm{age}\gtrsim 11.5$ Gyr).
@@ 3 The Auriga simulation suite
Figure 4: Circularity vs. bar strength for mono-age populations in Au18, 26 and 23. The colour of each dot represents the average age of the mono-age population. Each mono-age population are binned with 0.5 Gyr age segments. The black dashed lines label the bar strength threshold at $A_{2}=0.25$. The dot with black edges labels the population that born at the moment of bar formation in each Auriga galaxy.
@@ 4.1 Connection between the disc formation and bar signatures
Fig. 3 shows the results of the analysis of the kinematics of different mono-age populations in the inner galaxy of Au18 and Au26 using the snapshot at $\tilde{z}=0$. The disc formation history of Au18 and Au26 are revealed using the column-normalised (see e.g. Belokurov et al., 2020, for a similar representation of the Galaxy’s kinematics) age-circularity plot as shown in the top row of Fig. 3, in which the $\mathrm{circularity}=L_{z}/L_{\mathrm{circ}}$ is defined as the ratio of the angular-momentum in the $z$-direction to the angular-momentum of a circular orbit of equivalent energy. The orbits of the oldest stars in both galaxies are halo-like with a low and wide distribution of circularity. They have a slight prograde rotation as the mean circularity is greater than 0, similar to the results in other simulations (McCluskey et al., 2023; Chandra et al., 2023) and in observations (Chandra et al., 2023; Zhang et al., 2024b; Ardern-Arentsen et al., 2024). The trend of circularity as a function of age is shown by the red-solid line. Both Au18 and Au26 have a rapid spin-up with a short timescale of $\lesssim 0.5$ Gyr at a similar look-back time, $\sim 11.7$ Gyr and $9.2$ Gyr ago, respectively. We denote these as the spin-up time, $t_{\mathrm{spin-up}}$, for Au18 and Au26. The circularity of stars born after the spin-up stays at a high value, marking the existence of a long-lived disc.
@@ 4.3 Bar signature in the $v_{\ell}$ map
Figure 6: An illustration on the effects of the galactic bar on the $v_{\ell}$ field of an N-body simulated galaxy. Top left: the $v_{\ell}$ field of the mock galaxy. Top right: the $v_{\ell}$ field of the corresponding disc model we constructed for this mock galaxy. Bottom: the residual, $\Delta v_{\ell}$, map, in which we find an inclined, dipole-like signature that corresponds to the galactic bar. The black dashed ellipse is used to denote the size and orientation of the galactic bar in the N-body simulated galaxy
@@ 5 Spin-up epoch of the Milky Way
We split our Mira variable sample into different period bins, which is analogous to binning them by age thanks to the period-age relation of Mira variables (Wyatt & Cahn, 1983; Feast, 2009; Grady et al., 2019; Zhang & Sanders, 2023). The behaviour of Mira candidates in configuration space is shown in the first and second rows in Fig. 8. From the spatial distribution in the $(x,y)$ plane, an inclined, bar-like structure is seen in the period bins with $\mathrm{period}>190$ days. A bar-like overdensity is also found in populations with a period of 150-190 days but instead it clearly aligns with the $x$-axis, similar to the results in Grady et al. (2020). Distance uncertainty can bias the inclination angle of the Galactic bar by smearing the bar stars along the line-of-sight direction, and hence, shearing the bar so that it becomes more aligned with the Sun-Galactic centre (GC) line (Hey et al., 2023; Zhang et al., 2024a). However, to make the bar completely aligned with the Sun-GC line, distance uncertainty should be $25\%-30\%$, whereas for our sample, the distance uncertainty is only $\sim 10\%-15\%$ as shown in Fig. 2. It is thus clear that large distance uncertainty can also create an elongated, bar-like overdensity that is perfectly aligned with the Sun-GC line without there even being a real galactic bar. A detailed discussion of this matter can be found in Appendix A. In brief, as the heliocentric distance uncertainty displaces and smears stars along the line-of-sight direction, it stretches a spherical-like overdensity into a bar-like overdensity oriented along the line-of-sight. Therefore, the horizontal (in $x-y$ projection) bar-like signature we find in the Mira variables with a period of 150-190 days could be associated with a quasi-spherical overdensity in the inner Galaxy, for example, a spherical bulge or inner halo.
From the side-on density projection shown in the second row of Fig. 8, a larger scale height is found for Mira candidates with shorter periods (corresponding to the older population), which agrees with the results in Queiroz et al. (2023). Directly discussing the existence of the disc components from the spatial configuration of these populations is challenging as the age uncertainty affects the conclusion, as discussed in detail in Section 6.2. Hence, we avoid to over-interpreting the side-on projections of the Mira variables with a period between 150-190 days.
@@
The width of the period bin in Zhang & Sanders (2023) is generally smaller than the period bin adopted in this work, so the difference in the binning strategy has to be considered when comparing the disc model to the observation. For each of our period bins (which covers 2-3 period bins in Zhang & Sanders 2023) mock stars are generated from several disc models, for which the fractional contribution from each disc model is assigned according to the period distribution in our period bin. To fairly compare the $v_{\ell}$ map of the disc model and the observation, we scatter mock generated disc stars with $12\%$ distance uncertainty. Then, we match the distribution of the mock stars to the observed spatial distribution of our sample by finding the closest matched mock star for each observed star in the configuration space. This routine is repeated five times to minimise the effect of Poisson uncertainty on the disc model. We then subtract the $v_{\ell}$ map of the disc model from the observation. In the lower row of Fig. 8, we show the residual $v_{\ell}$ map in terms of $\Delta v_{\ell}/\sigma_{v_{\ell}}$ for each period bin, where $\sigma_{v_{\ell}}$ is the Poisson error for each $(x,y)$ pixel. A strong dipole signal that is inclined like the Galactic bar is seen in the long period (young) population as expected due to the Galactic bar. By visual inspection, the dipole signal gradually becomes weaker for a shorter period (older) population and no dipole-like signature is noticeable in the 150-190 period Mira variables.
@@
where $\tau_{0}$ and $P_{s}$ are the associated parameters. Zhang & Sanders (2023) fitted Eq. 6 and found $(\tau_{0},P_{s})$ = (13.7 Gyr, 401 days) when considering Mira kinematics only or $(\tau_{0},P_{s})$ = (14.7 Gyr, 308 days) when information of Mira variables in globular clusters is included. Zhang & Sanders (2023) found an age scatter of $11\%$ at a fixed period based on the difference in the Mira ages based on kinematics and on globular cluster ages. In this work, we adopt the $11\%$ scatter as the intrinsic dispersion in the period-age relation. However, the true intrinsic dispersion may require further calibration as theoretical models predict a significant intrinsic scatter (Trabucchi et al., 2019). Mira variables with a period of 190 days have ages corresponding to $10.5\pm 1.2$ Gyr and $9.1\pm 1.0$ Gyr for $(\tau_{0},P_{s})$ = (14.7 Gyr, 308 days) and $(\tau_{0},P_{s})$ = (13.7 Gyr, 401 days), respectively. Therefore, the age of the Milky Way disc is likely to be younger than $\sim 12$ Gyr (10.5 + 1.2 Gyr).

### v3

# Deciphering the Milky Way disc formation time encrypted in the bar chrono-kinematics

@@ Abstract
pubyear: 2024pagerange: Deciphering the Milky Way disc formation time encrypted in the bar chrono-kinematicsA
@@ 1 Introduction
A disc is a typical component of many galaxies in the Universe. Disc formation and evolution in galaxies are studied with high-redshift observations of morphology (e.g. Nelson et al., 2023; Kartaltepe et al., 2023; Robertson et al., 2023) and gas kinematics (e.g. Wisnioski et al., 2019; Übler et al., 2019; de Graaff et al., 2024), as well as with cosmological simulations (e.g. Stern et al., 2021; Hopkins et al., 2023; Semenov et al., 2024; Dillamore et al., 2024a). In the Milky Way, because we can resolve individual stars and measure their complete 6D phase space positions, our Galaxy is an excellent laboratory to study disc formation and evolution (e.g. Carollo et al., 2019; Sestito et al., 2020; Di Matteo et al., 2020; Chandra et al., 2023; Zhang et al., 2024b; Nepal et al., 2024a; Gallart et al., 2024).
@@
The formation time of the Milky Way disc can be studied both using metallicity as a clock of Galactic evolution (Norris et al., 1985; Sestito et al., 2020; Di Matteo et al., 2020; Belokurov & Kravtsov, 2022; Chandra et al., 2023; Zhang et al., 2024b; Fernández-Alvar et al., 2024; Viswanathan et al., 2024) as well as using stellar ages directly (Miglio et al., 2021; Rix et al., 2022; Xiang & Rix, 2022; Queiroz et al., 2023; Nepal et al., 2024a). The metallicity distribution of the Galactic disc has received much attention in the past decades, especially at the metal-poor regime, which is potentially linked to the epochs of its early formation. Observations invoking the existence of a metal-weak thick disc are discussed periodically (e.g. Norris et al., 1985; Morrison et al., 1990; Chiba & Beers, 2000; Carollo et al., 2010, 2019; An & Beers, 2020; Di Matteo et al., 2020).
@@
Recently, very metal-poor (VMP) stars, which are expected to be the oldest stars in the Milky Way, with disc-like orbits discovered by Sestito et al. (2019, 2020) have revived the discussion. Although the disc-like VMP stars could have multiple origins  (see Sestito et al., 2021), the hypothesis that they come from an early proto-disc has attracted the most attention (e.g. Di Matteo et al., 2020; Mardini et al., 2022; Bellazzini et al., 2024). Using APOGEE data, Belokurov & Kravtsov (2022) created a pure sample of in-situ MW stars and revealed that the net azimuthal velocity of the Galaxy increases abruptly across a short metallicity range around $\mathrm{[M/H]}\sim-1.3$. Comparing to a variety of numerical models of galaxy formation, Belokurov & Kravtsov (2022) conclude that the fast observed spin-up of the Milky Way marks the emergence of the stable and dominant Galactic stellar disc. Note that while the metallicity scale of the Galactic spin-up is not very low, the corresponding lookback time is impressively large, corresponding to redshifts above $z>3$. Further support for this hypothesis has been found in independent studies of other simulation suites (e.g. Semenov et al., 2024; Dillamore et al., 2024a; Chandra et al., 2023). Indeed, as these studies demostrate, our Galaxy was likely one of the first to form a stable stellar disc amongst objects of similar halo mass. With metallicities from the Gaia BP/RP (XP) spectra (e.g. Andrae et al., 2023), we can access an unprecedentedly large sample of metal-poor stars. Recent analyses of such Gaia XP-based samples bolster the idea that the Milky Way disc formed around $\mathrm{[M/H]}\sim-1.5$ to $-1$ (Chandra et al., 2023; Zhang et al., 2024b; Kane et al., 2024). Other origins for the VMP stars on disc-like orbits have also been discussed. These can form in the proto-Galaxy  (Sestito et al., 2021; Horta et al., 2024; Zhang et al., 2024b), or can be trapped by a decelerating bar (Dillamore et al., 2023; Li et al., 2023; Yuan et al., 2023).
@@
Although the stellar metallicity is usually measured more accurately than the age, the metallicity evolution also depends on the star formation history (sometimes leading to non-monotonic age-metallicity relation), which complicates the interpretation of results. Accurate age measurements of disc stars can directly pin down the formation epoch of the Milky Way disc, but this is hard, and very few robust age estimates exist for low-metallicity stars. Miglio et al. (2021) assessed the asteroseismological age of stars with light curves measured by the Kepler satellite and found the mean age of stars in the thick disc after deconvolving the age uncertainty is consistent with $\sim 11$ Gyr. Similarly, Queiroz et al. (2023) used the stellar ages from isochrone fitting and found the same age for the thick disc. Recently, Nepal et al. (2024a) found disc populations of stars older than $13$ Gyr using stellar parameters from the Gaia RVS+XP (Guiglion et al., 2024), which hints that the Galactic disc may have started to form within $1$ Gyr after the Big Bang. However, reconciling the early disc formation scenario with the cosmological simulations is challenging. In numerical models, the morphology of a galaxy appears to change quickly from spheroidal to disc-like when the galaxy reaches a critical mass threshold (Hopkins et al., 2023; Dillamore et al., 2024a). Taking the diversity of the assembly histories into account, it appears unusual for a Milky Way-like galaxy to reach such a mass threshold much earlier than redshift $\sim 3$ (Belokurov & Kravtsov, 2022).
@@
Previous measurements of the disc formation epoch use the kinematics of the solar vicinity or globular clusters. In this work, we provide a new independent method to measure the disc formation epoch that instead uses the chrono-kinematics of the Galactic bar. The Galactic bar dominates the inner Galaxy and is identified from the stellar density distribution at the centre (Nakada et al., 1991; Stanek et al., 1997; Wegg & Gerhard, 2013). Other than the spatial overdensity caused by the Galactic bar, it also induces distinct kinematic features in the inner Galaxy, which are useful when tracing the bar signatures. Due to the streaming motion of bar-supporting stars and their elongated orbits, the kinematics of the Galactic bar is characterised by the quadrupole radial velocity pattern (Bovy et al., 2019; Queiroz et al., 2021; Hey et al., 2023; Vislosky et al., 2024; Liao et al., 2024; Zhang et al., 2024a) and the bi-symmetric pattern in the fractional radial velocity map (Zhang et al., 2024a).
@@
The structure and formation of the bar has been explored intensively with N-body simulations (e.g. Athanassoula & Misiriotis, 2002; Fragkoudi et al., 2017; Fujii et al., 2018; Ghosh et al., 2024; Jiménez-Arranz et al., 2024), and recently, with the Auriga cosmological simulation suite (Grand et al., 2017; Fragkoudi et al., 2020; Merrow et al., 2024; Fragkoudi et al., 2024). Several factors can impact the bar formation in a galaxy, such as disc kinematics (e.g. Toomre, 1964; Athanassoula & Sellwood, 1986), disc mass fraction and its assembly (e.g. Fujii et al., 2018; Bland-Hawthorn et al., 2023; Khoperskov et al., 2024), and gas fraction (e.g. Athanassoula et al., 2013), but the exact conditions are still complicated (Romeo et al., 2023). Many studies used N-body simulations to investigate the response of different stellar populations to the Galactic bar (e.g. classical bulge: Saha et al. 2012, 2016, the stellar halo: Pérez-Villegas et al. 2017, and the Galactic disc: Debattista et al. 2017; Fragkoudi et al. 2017). Debattista et al. (2017) demonstrated that cold populations form a strong, thin, peanut-shaped bar, while hot populations form a weak, thick, boxy-shaped bar. As kinematically colder population respond to the Galactic bar more efficiently, the cold galactic disc provides the stars for the bar formation (Toomre, 1964; Fragkoudi et al., 2017; Debattista et al., 2017; Boin et al., 2024). Moreover, early interactions with large satellites can seed bar formation. For example, Merrow et al. (2024) show that an accretion event similar to the GS/E merger (Belokurov et al., 2018; Helmi et al., 2018) in the MW can be responsible for the bar formation, as long as the GS/E progenitor is not too massive.
@@
Long-period variables (LPV) are bright stars that follow tight period-luminosity relations (Rau et al., 2019; Sanders, 2023) and hence, are often used to trace the structure and kinematics of the inner Galaxy (e.g. Catchpole et al., 2016; Grady et al., 2020; Semczuk et al., 2022; Sanders et al., 2022; Hey et al., 2023; Zhang et al., 2024a). Mira variables, a subcategory of LPVs, are high-amplitude thermally pulsating asymptotic giant branch stars. Their distances can be calibrated via the period-luminosity relation of O-rich Mira variables in the Milky Way (Sanders, 2023) and can have astrometry from Gaia DR3. Mira variables also exhibit a period-age relation: Miras with shorter periods have older ages (Feast & Whitelock, 2000; Grady et al., 2019; Trabucchi & Mowlavi, 2022; Zhang & Sanders, 2023). Zhang & Sanders (2023) re-calibrated the period-age relation of O-rich Mira variables to ensure consistency with the ages of Miras in globular clusters. Mira variables have been used to trace the chronological evolution of the Milky Way structures (e.g. Catchpole et al., 2016; Grady et al., 2020; Semczuk et al., 2022), and recently, Sanders et al. (2024) used the Mira variables in the nuclear stellar disc to constrain the bar formation epoch of the Milky Way.
@@ 2.1 O-rich Mira variable selection
Although Mira variables with a period shorter than $150$ days exist, we exclude LPV candidates with $\mathrm{period}<150$ days to remove potential short-period(SP)-red stars contamination. (e.g., Feast & Whitelock, 2000). Kinematically, the SP-red stars behave similarly to longer-period Mira variables, which have a lower velocity dispersion. This indicates that rather than being older these stars are less dynamically heated and so have smaller ages. A related result is found in Zhang & Sanders (2023), in which the kinematics of the Mira sample with a period of 80 to 150 days is kinematically colder than expected. We omit the contaminating stars, many of which are likely SP-red objects from our analysis by neglecting Mira candidates with a period shorter than 150 days, which also removes the oldest Mira candidates ($\mathrm{age}\gtrsim 11.5$ Gyr).
@@ 3 The Auriga simulation suite
Figure 4: Circularity vs. bar strength for mono-age populations in Au18, 26 and 23. The colour of each dot represents the average age of the mono-age population. Each mono-age population are binned with 0.5 Gyr age segments. The black dashed lines label the bar strength threshold at $A_{2}=0.25$. The square/triangle with black edges labels the circularity and bar strength for the population that born at the moment of bar formation/spin-up in each Auriga galaxy.
@@ 4.1 Connection between the disc formation and bar signatures
Fig. 3 shows the results of the analysis of the kinematics of different mono-age populations in the inner galaxy of Au18 and Au26 using the snapshot at $\tilde{z}=0$. The top panels are made with all stellar particles in Au18 and 26, while the lower set of panels focuses on particles in the inner galaxy with $|x|<5$ and $|y|<5$ kpc. The disc formation history of Au18 and Au26 are revealed using the column-normalised (see e.g. Belokurov et al., 2020, for a similar representation of the Galaxy’s kinematics) age-circularity plot as shown in the top row of Fig. 3, in which the $\mathrm{circularity}=L_{z}/L_{\mathrm{circ}}$ is defined as the ratio of the angular-momentum in the $z$-direction to the angular-momentum of a circular orbit of equivalent energy. The orbits of the oldest stars in both galaxies are halo-like with a low and wide distribution of circularity. They have a slight prograde rotation as the mean circularity is greater than 0, similar to the results in other simulations (McCluskey et al., 2023; Chandra et al., 2023) and in observations (Chandra et al., 2023; Zhang et al., 2024b; Ardern-Arentsen et al., 2024). The trend of circularity as a function of age is shown by the red-solid line. Both Au18 and Au26 have a rapid spin-up with a short timescale of $\lesssim 0.5$ Gyr at a similar look-back time, $\sim 11.7$ Gyr and $9.2$ Gyr ago, respectively. We denote these as the spin-up time, $t_{\mathrm{spin-up}}$, for Au18 and Au26. The circularity of stars born after the spin-up stays at a high value, marking the existence of a long-lived disc.
@@ 4.3 Bar signature in the $v_{\ell}$ map
Figure 6: An illustration on the effects of the galactic bar on the $v_{\ell}$ field of an N-body simulated galaxy. Top left: the $v_{\ell}$ field of the mock galaxy. Top right: the $v_{\ell}$ field of the corresponding disc model we constructed for this mock galaxy. Bottom: the residual, $\Delta v_{\ell}$, map, in which we find an inclined, dipole-like signature that corresponds to the galactic bar. The black contours show the surface density distribution of stellar particles in this galaxy.
@@ 5 Spin-up epoch of the Milky Way
We split our Mira variable sample into different period bins, which is analogous to binning them by age thanks to the period-age relation of Mira variables (Wyatt & Cahn, 1983; Feast, 2009; Grady et al., 2019; Zhang & Sanders, 2023). The behaviour of Mira candidates in configuration space is shown in the first and second rows in Fig. 8. From the spatial distribution in the $(x,y)$ plane, an inclined, bar-like structure is seen in the period bins with $\mathrm{period}>190$ days. A bar-like overdensity is also found in populations with a period of 150-190 days but instead it clearly aligns with the $x$-axis, similar to the results in Grady et al. (2020). Distance uncertainty can bias the inclination angle of the Galactic bar by smearing the bar stars along the line-of-sight direction, and hence, shearing the bar so that it becomes more aligned with the Sun-Galactic centre (GC) line (Hey et al., 2023; Zhang et al., 2024a). However, to make the bar completely aligned with the Sun-GC line, distance uncertainty should be $25\%-30\%$, whereas for our sample, the distance uncertainty is only $\sim 10\%-15\%$ as shown in Fig. 2. A distance uncertainty on the order of $10\%$ can also create an elongated, bar-like overdensity that is aligned with the Sun-GC line from an approximately spherical distribution. We demonstrate this in Appendix A with details. In brief, as the heliocentric distance uncertainty displaces and smears stars along the line-of-sight direction, it stretches a spherical-like overdensity into a bar-like overdensity oriented along the line-of-sight. Therefore, the horizontal (in $x-y$ projection) bar-like signature we find in the Mira variables with a period of 150-190 days could be associated with a quasi-spherical overdensity in the inner Galaxy, for example, a spherical bulge or inner halo.
From the side-on density projection shown in the second row of Fig. 8, a larger scale height is found for Mira candidates with shorter periods (corresponding to the older population), which agrees with the results in Queiroz et al. (2023). Directly discussing the existence of the disc components from the spatial configuration of these populations is challenging as the age uncertainty affects the conclusion, as discussed in detail in Section 6.2. Hence, we avoid over-interpreting the side-on projections of the Mira variables with a period between 150-190 days.
@@
The width of the period bin in Zhang & Sanders (2023) is generally smaller than the period bin adopted in this work, so the difference in the binning strategy has to be considered when comparing the disc model to the observation. For each of our period bins (which covers 2-3 period bins in Zhang & Sanders 2023) mock stars are generated from several disc models, for which the fractional contribution from each disc model is assigned according to the period distribution in our period bin. The mock stars are generated with the fitted quasi-isothermal distribution function parameters in Zhang & Sanders (2023) using the Agama Sampling routine (see details in Vasiliev 2019). To fairly compare the $v_{\ell}$ map of the disc model and the observation, we scatter mock generated disc stars with $12\%$ distance uncertainty. Then, we match the distribution of the mock stars to the observed spatial distribution of our sample by finding the closest matched mock star for each observed star in the configuration space. This routine is repeated five times to minimise the effect of Poisson uncertainty on the disc model. We then subtract the $v_{\ell}$ map of the disc model from the observation. In the lower row of Fig. 8, we show the residual $v_{\ell}$ map in terms of $\Delta v_{\ell}/\sigma_{v_{\ell}}$ for each period bin, where $\sigma_{v_{\ell}}$ is the Poisson error for each $(x,y)$ pixel. A strong dipole signal that is inclined like the Galactic bar is seen in the long period (young) population as expected due to the Galactic bar. By visual inspection, the dipole signal gradually becomes weaker for a shorter period (older) population and no dipole-like signature is noticeable in the 150-190 period Mira variables.
@@
where $\tau_{0}$ and $P_{s}$ are the associated parameters. Zhang & Sanders (2023) fitted Eq. 6 and found $(\tau_{0},P_{s})$ = (13.7 Gyr, 401 days) when considering Mira kinematics only or $(\tau_{0},P_{s})$ = (14.7 Gyr, 308 days) when information of Mira variables in globular clusters is included. Zhang & Sanders (2023) found an age scatter of $11\%$ at a fixed period based on the difference in the Mira ages based on kinematics and on globular cluster ages. In this work, we adopt the $11\%$ scatter as the intrinsic dispersion in the period-age relation. However, the true intrinsic dispersion may require further calibration as theoretical models predict a significant intrinsic scatter (Trabucchi et al., 2019). Mira variables with a period of 190 days have ages corresponding to $10.5\pm 1.2$ Gyr and $9.1\pm 1.0$ Gyr for $(\tau_{0},P_{s})$ = (14.7 Gyr, 308 days) and $(\tau_{0},P_{s})$ = (13.7 Gyr, 401 days), respectively. Therefore, the age of the Milky Way disc is likely to be younger than $\sim 12$ Gyr (10.5 + 1.2 Gyr). This result is also consistent with the observed age distribution of the Galactic bulge in Bensby et al. (2013).